1. A project will deliver $90 now, -$10 in a fortnight's time, and $30 in two fortnight's time. The appropriate interest/discount rate is j26 = 8.060% p.a. The Present Value (i.e period 0) of the ENTIRE cashflow delivered is: 1) $90.00; 2) $110.53; 3) $106.44; 4) $109.85
Answer #1

Solution:

Discount rate per annum = 8.060%

Discount rate fortnightly = 8.060/52*2= 0.31%

Present value of cash flow = 90/(1.0031)^0 - 10/(1.0031)^1 + 30/(1.0031)^2 = $109.85

Option 4 is correct i.e. $109.85